LOG-IN
Displaying reports 3241-3260 of 3360.Go to page Start 159 160 161 162 163 164 165 166 167 End
KAGRA MIF (General)
Print this report.
YoichiAso - 06:05, Thursday 16 July 2015 (130)Get code to link to this report
Comment to Preparing the instrument control racks for placing in the mine (Click here to view original report: 129)

I checked the temperature inside the rack after about 24 hours. It was 23.9C. The temperature raise is too small. I will look for a way to generate more heat (heater) in a controlled way (slidac ?).

KAGRA MIF (General)
Print this report.
YoichiAso - 16:20, Tuesday 14 July 2015 (129)Get code to link to this report
Preparing the instrument control racks for placing in the mine

At some point, we have to put the instrument control system racks in the humid environment of the tunnel.
I borrowed a plastic cover from DGS and put it on one of the racks. Then I turned on the system. I also put a temperature meter.
I will leave it like this until tomorrow to see if the temperature goes up high enough (above 30C).

If the temperature is high, we can expect the local humidity to drop accordingly. So we can move the rack to the field.

Right now, the wrapped rack is sitting in front of the shutter.


 

Images attached to this report
129_20150714091922_20150714132718.jpg
Comments related to this report
YoichiAso - 06:05, Thursday 16 July 2015 (130)

I checked the temperature inside the rack after about 24 hours. It was 23.9C. The temperature raise is too small. I will look for a way to generate more heat (heater) in a controlled way (slidac ?).

KAGRA VIS (OSEM)
Print this report.
TomotadaAkutsu - 20:24, Monday 13 July 2015 (128)Get code to link to this report
Comment to Hmm... (Click here to view original report: 126)

In order to see how the intensity noises of the LED (TSTS7100, not OP232) will change, I slightly modify the Okt driver. A trimmer (CT-94EW500-ohm) is attached to the driver so that the LED current can be variable. The attached graphs are the measurement results. According to the results, when the DC powers is over a certain level (in this case it is around 4Vdc... note that which is caught by the OSEM prototype setup, not the naked direct input power to the PD from the LED) , the intensity noises are not proportional to the DC intensities.

The followings can be said:

(1) because we don't want to input such a large current into vacuum, and the lowest usable DC current should be supplied to the LED (I still did not measure the trimmer's resistance). In terms of this, the DC current should be tuned to make the PD output to 1Vdc~2Vdc.

(2) If you are worrying about the other noises (scattering light of 1064nm, analog-to-digital noise, and ...) would cover the OSEM signals, the calibration factor (or sensitivity) of the OSEM should be kept high, meaning larger DC current to LED would be better. Then the maximum PD output would be  (this is only my recommendation/suggestion) around 4Vdc.

 

My recommendation:

I like to setup the PD output to 1.5Vdc, which would be sufficient.

Images attached to this comment
128_20150713132318_plot20150713pub.png 128_20150713132323_plot20150713compare1.png 128_20150713132331_plot20150713compare2.png 128_20150713132336_plot20150713compare3.png 128_20150713132342_plot20150713compare4.png 128_20150713132348_plot20150713compare5.png 128_20150713132353_plot20150713compare6.png 128_20150714025943_plot20150713compares.png
R&D (Cryogenic)
Print this report.
SimonZeidler - 14:32, Monday 13 July 2015 (127)Get code to link to this report
Cooling-down curve and warming-up

The cooling-down has been finished at Friday night. The final temperature is 4K, just as given also by the company.
The overall duration of cooling down with full vacuum is about 55 hours.

I started the warming-up today in the morning just by switching off the cryo-cooler.
The temperature curve is being recorded again.

Attached are a photograph of the temperature controller this morning and the cooling-curve as taken during the test.

 

Images attached to this report
127_20150713073143_150713103046.jpg 127_20150713073333_calibration.jpg
KAGRA VIS (OSEM)
Print this report.
TomotadaAkutsu - 20:38, Friday 10 July 2015 (126)Get code to link to this report
Hmm...

The attached graph shows noise level of two OSEM setups; with a satellite box (Sat) and with a test driver (Okt) made in NAOJ (design based on the Sat, but Akutsu chooses necessary/sufficient parts what he thought, and Okutomi made). Those two circuits supply different amount of DC current to the LED (TSTS7100). Exactly the same LED and PD combination is used for the measurements.

(0) The measurement systems's noise level (raw voltage noises) for both cases when the LEDs were turned off are in the same order of magnitude. (I don't mention the PD's noises here, and it will be mentioned in (1). I'm mentioning about the noise of measurement system with the range which is used when the PD circutis are measured. The second attached graphs are the raw voltage spectra (V/rtHz vs Freq), and showing this. S70908 (for the satellite box) measurement was done with 6.3mVpk range, while S70905 (for the Okt driver) was 7.95mVpk, and the ratio can explain the relation of both measurement noises) The difference shown are due to the difference between calibration factors (V/m) for the cases.

(1) Why are the PD noises (LED off) such different...?? For the Okt test driver, OP27G is used; for the satellite box, it seems OP2177. The transimpedances are the same (38.3k Ohm).

(2) The LED's intensity noises do not appear proportional to the DC powers. The displacement noises should be the same level if it is proportional.

(3) Additional Note: the satellite box's output of PD driver has a relatively large offset like 20mVdc even without any PDs are not connected, and which is not observed in the Okt test driver. I just add a 5.6k Ohm (by clips) parallely to the 38.3k Ohm and observed the offset decreases accordingly (38.3//5.6 = 4.88kOhm, and 38.3/4.88 = 7.84 times), and the same happens for the power spectrum of the intensity noise... so the offset would come from the opamp's current offset and current noise-like things. As was reported, some polarized capacitances are directed wrong, so some opamps would have malfunctinos now... I checked a OP2177's data sheet but cannot find descriptions on the current noise around low frequencies. The current offset is described but it is too low to explain what are happening here.

(4) Next I decied to use only the Okt test driver's PD circuit for evaluating what's going on the LED noise. With this driver, the larger (and non-proportionally large) noise of the LED driven by the satelite box is confirned again. The current to the LED would be now estimated to 88.3mA (the feedback loop makes 200//33 Ohms should be 2.5V)with the satellite box, and 8.3mA (the feedback loop makes 200+100 Ohms should be 2.5V) with the Okt test driver. This could explain, at least order, the rate why the output is 8.92Vdc for the satellite box and 740mVdc for the Okt. The intensity noise spectra are, however, not proportional to those DC voltages or current to the LEDs. Need to be more investigated; the goal would be to find the LED current setup in which the intensity noise would be slightly higher than the PD driver's noise; fortunately, I think the Okt circuit condition is very closed to this goal!

Anyway, generally speaking, I recommend to use a FET-input opamp for a transimpedance amplifier application like the PD driver circuit... as I did for the oplev QPD driver for KAGRA.

Images attached to this report
126_20150710133026_figure1.png 126_20150710165321_noises.png
Comments related to this report
TomotadaAkutsu - 20:24, Monday 13 July 2015 (128)

In order to see how the intensity noises of the LED (TSTS7100, not OP232) will change, I slightly modify the Okt driver. A trimmer (CT-94EW500-ohm) is attached to the driver so that the LED current can be variable. The attached graphs are the measurement results. According to the results, when the DC powers is over a certain level (in this case it is around 4Vdc... note that which is caught by the OSEM prototype setup, not the naked direct input power to the PD from the LED) , the intensity noises are not proportional to the DC intensities.

The followings can be said:

(1) because we don't want to input such a large current into vacuum, and the lowest usable DC current should be supplied to the LED (I still did not measure the trimmer's resistance). In terms of this, the DC current should be tuned to make the PD output to 1Vdc~2Vdc.

(2) If you are worrying about the other noises (scattering light of 1064nm, analog-to-digital noise, and ...) would cover the OSEM signals, the calibration factor (or sensitivity) of the OSEM should be kept high, meaning larger DC current to LED would be better. Then the maximum PD output would be  (this is only my recommendation/suggestion) around 4Vdc.

 

My recommendation:

I like to setup the PD output to 1.5Vdc, which would be sufficient.

Images attached to this comment
128_20150713132318_plot20150713pub.png 128_20150713132323_plot20150713compare1.png 128_20150713132331_plot20150713compare2.png 128_20150713132336_plot20150713compare3.png 128_20150713132342_plot20150713compare4.png 128_20150713132348_plot20150713compare5.png 128_20150713132353_plot20150713compare6.png 128_20150714025943_plot20150713compares.png
R&D (Cryogenic)
Print this report.
SimonZeidler - 18:06, Wednesday 08 July 2015 (125)Get code to link to this report
First cycle of cooling

The first cycle of cooling the tank has begun today.
After having some trouble with the calibration of the temperature controller and a sudden (and still unexplained) problem with the pressure meter, we are now cooling the tank.

The rate of cooling is now approximately 0.07 K/min at an inside pressure of 2.2*10^(-4) Pa.
With this rate, the cooling down to 10 K will last at least 2,5 days (even more is expected when the temperature falls below 70 K and freeze-out of the remaining gas appears).

I think of using He gas in the future that could be injected into the tank for increasing the effectivity of the cooling process.

 

Attached is a photograph of the temperature and the pressure controller at the beginning of the cooling cycle, right after the cool-down has started.

Images attached to this report
125_20150708110752_150708132415.jpg
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 14:55, Tuesday 07 July 2015 (124)Get code to link to this report
Power dependence of absorption

I measure the surface absorption peak of LMA sample #15033 at different powers.

The first plot shows the scans along Z axis

The second plot shows AC signal peak at the surface as a function of the pump power.

The third plot shows the proportionality as a function of the pump power

Images attached to this report
124_20150707075330_pumppower1.jpg 124_20150707075338_pumppower2.jpg 124_20150707075345_pumppower3.jpg
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 02:40, Tuesday 07 July 2015 (123)Get code to link to this report
LMA sample 15033

Last friday (3rd July) I cleaned both surfaces of LMA coated sample #15033 (abs=12.8ppm) as shown in the first picture using the first contact polymeric optics cleaner.

I measured the absorption of the sample along Z axis. I made scans of 10mm in order to see both surfaces. The sample thickness is 6mm.

I made a scan (blue line in the second figure) and then I flipped the sample in order to see also the other surface (black line) and I notices a strange peak, then I moved the sample 0.1mm along Y axis and I repeated the scan (red line). The black line shows an absorption that has a not usual pattern. I guess it is because of some scrach on the surface, because just 0.1mm apart it doesn't show.

This monday (6th July) I made again the measurement of sample and of flipped sample, (third picture). It shows a variation among the two surfaces of a factor of 2. Maybe it is just a fluctuation of the absorption value because flipping the sample we never go back to the same X,Y position.

Images attached to this report
123_20150706192918_rimg0002.jpg 123_20150706192937_plotlma15033friday.jpg 123_20150706194202_plotlma15033monday.jpg
Comments related to this report
ManuelMarchio - 12:36, Saturday 18 July 2015 (135)

I wanted to understand the big difference between the absorption peak in the configuration where the pump is entering in the sample by the absorbing surface and the other configuration where the pump is entering the sample by the not absorbing surface (flipped sample). So I made again a calibration and also a calibration with the flipped surface reference sample and I found a calibration factor R=17 1/W (flipped) and R=13 1/W (not flipped). For the not flipped sample I made a calibration history in this post and I found a mean calibration factor R=11 1/W with fluctuatoins of 10%. I don't understand why R changes so much. 

If we calculate the LMA15033 absorption with those R factors we get 22ppm for the flipped surface and 14ppm for the not flipped surface.

R&D (Cryogenic)
Print this report.
SimonZeidler - 14:16, Friday 03 July 2015 (122)Get code to link to this report
Test of Vacuum System

We have finished the third cycle of pumping and observing the pressure after closing the gate valve on the cryostat in the ATC.

The results, compared with the other two cycles and the initial measurement, can be seen in the attached pictures.
Obviously, for the pumping, we can not get the vacuum that has been reached in the initial measurements but we could confirm the data from the second cycle with a final pressure of approx. 6*10^(-4) Pa after 60000s of pumping. The initial cycle reached approx. 4*10^(-4) Pa.
Again, the vacuum is by far better than in the first cycle after the cryostats repair.

The rising of the pressure after closing the gate valve follows quite exactly the development of the initial cycle. No important differences could be observed.

We also checked the connections for the temperature controller and are ready to braze the cabels on the respective connectors (will take only 30min, or so).
From our understanding, we are ready for doing the test of the cryogenic system and will start it on Monday, probably.

Non-image files attached to this report
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 11:24, Thursday 02 July 2015 (121)Get code to link to this report
Measurement of a low absorption sample of LMA

I put the surface calibration sample in the big sample holder and I found the position of the sample surface along the Z axis is about 8mm different with respect to the little reference sample holder. This just means the scan should start from Z0=12mm instead of Z0=20mm.

I put the sample of LMA  in the holder and I attached the holder to the translation stage, I acquired 1hour of noise and it turns out to be 18microV std with no spiky noise. This is means the presence of the sample holder doesn't increase the noise.

I acquired the scan of the coated sample they gave me at LMA (sample 15033, they measured abs=12.8ppm @1064). Pump power = 171mW (LD current = 1.2A)

The first image shows the scan along Z axis of the sample. We can see two peaks and this is not the coating absorption patter we see in the calibration. The reason is that we are looking at the surface which doesn't have a HR coating

The second image shows the scan along Z axis of the flipped sample, in order to see the right HR coated surface. It has a pattern very similar to the calibration scan, a high peak at the surface and the two smaller peaks of diffraction effects.

The absorption calculated from this measurements and the calibration factor with the reference sample we have (surf_abs = 22%) gives 107ppm, almost a factor of 8 higher than the nominal value (12.8ppm). We are understanding the meaning of this. We will check the calibration factor, how the software filtering works and how clean is the surface of the sample.

The formula is Abs=AC/(DC*Power*R) where R is the calibration factor, it has a value of 10+/-1 and it is calculated with the surface reference sample calibration (abs=22%).

Images attached to this report
121_20150701164045_lma15033p171mwfilter.jpg 121_20150701164110_lma15033p171mwfilterswap.jpg 121_20150702101330_rimg0003768x1024.jpg
KAGRA VIS (OSEM)
Print this report.
TomotadaAkutsu - 15:01, Tuesday 30 June 2015 (120)Get code to link to this report
OSEM assembly procedure

I wrote an OSEM assembly procedure; please have a look!

http://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=3762

KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 17:14, Monday 29 June 2015 (119)Get code to link to this report
Comment to Strange AC noise (Click here to view original report: 117)

Today (Monday 29th) I acquired an hour of noise in the exactly same condition and I found 18microV of noise and no spiky noise at all. The only thing that changed since last time is that the system has been at rest for the weekend, about 48 hours.

KAGRA VIS (OSEM)
Print this report.
TomotadaAkutsu - 23:13, Friday 26 June 2015 (118)Get code to link to this report
Comment to OSEM LED screening preparation (Click here to view original report: 114)

[Abstract]

For measuring the LED intensity noise, the large DC voltage spoils the dynamic range of the FFT analyzer. The AC coupling could be usable to overcome the situation, so the difference between the AC and DC coupling inputs of 35670A's are measured. We should include the fit function for the calibration of measured data with the AC couping input.

[Measurements]

Measuring an input signal (or noise) with Ch1 and 2 of 35670A simultaneously and make it calculate the ratio Ch2/Ch1. A signal input is divided by a T-shape stuff and input into each channel.

  • (a) Freq. response of Ch2/CH1 are measured with a random noise input created by the 35670A itself.
  • (b) Swept sine mode to measure the same thing.

For the both cases, Ch2 is set to AC float and Ch1 to DC float. (I've tried other combination for just a confirmation.)

[Results]

The results of (a) and (b) are consistent, and  the fit transfer function is approximated by

f(x) = (i*x/x0)/(1+i*x/x0)

and

 x0 -> 0.485(5),

where I just assume f(x) -> 1 as x -> infinity.

You can use this function for the calbration when you use this 35670A with the AC coupling input (CH2).

Note: I'd like to avoid SR560s for the OSEM screening setup, as it appears to have large noise comparing to what I want to measure in the low frequency region, and each SR560 looks to have different noise level one by one... maybe our SR560s are getting old since they had been bought in 1998; one of them even sometimes has a jump in the DC offset during the measurement.

Images attached to this comment
118_20150626164136_bode20150626.png
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 20:14, Friday 26 June 2015 (117)Get code to link to this report
Strange AC noise

In order to check if the spiky noise decreases after placing the clean booth,  long time acquisition of AC noise was performed (1 hour).

I couldn't find spikes, but we see a strange beaviour: the noise increase with time to a very high level (1mV std). Then another acquisition of 5 minutes was performed and it can be seen the noise suddenly decrease alone at a certain point. I'm trying to find a reason to this strange behaviour. The booth fan are ON, I think they never stop alone. Maybe some fan of the air conditioning or other noisy devices, I don't know.

------------------------------------------------------------------------------------------------------------------------------------

I used a vacuum cleaner to clean the floor inside the clean booth and around it. The floor is not smooth, it's not easy to remove all the dust. 

Images attached to this report
117_20150626125842_probeon1hour.jpg 117_20150626125850_probeon5min.jpg
Comments related to this report
ManuelMarchio - 17:14, Monday 29 June 2015 (119)

Today (Monday 29th) I acquired an hour of noise in the exactly same condition and I found 18microV of noise and no spiky noise at all. The only thing that changed since last time is that the system has been at rest for the weekend, about 48 hours.

KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 18:03, Thursday 25 June 2015 (116)Get code to link to this report
Clean booth over the optical table

Members: Manuel, D.Tatsumi,

we clean up the optical table, remove unnecessary things, clean all cables, stick labels on cables, clean the floor.

We moved the clean booth from the upper floor to the optical table downstairs. And we moved the wheel desk from downstair to upstair.

Images attached to this report
116_20150625110041_img0130.jpg 116_20150625110050_img0133.jpg
KAGRA VIS (OSEM)
Print this report.
TomotadaAkutsu - 00:53, Thursday 25 June 2015 (114)Get code to link to this report
OSEM LED screening preparation

LEDs of OSEMs should be screened by measuring their intensity noise beforehand, so a setup for this business is prepared.

It is a quite simple setup holding a LED and PD closely with Thorlabs jigs (see pic).

For the driver for the LED and PD, today I just use the Okutomi-kun's circuit (made based on the actual OSEM satellite box design, but a little bit modified).

  • PD: Hamamatsu S1223-01
  • LED: OP232 (I'd like to change this to TSTS7100...)

When the LED is turned on, the DC ouput of the PD driver is 1.5 V; the transimpedance is 38.3k Ohm, so the photocurrent would be about 40uA. According to the specification sheet, the PD would have about 0.55 A/W at the wavelenth (OP232's light is 890nm), so the corresponding light flux collected by the PD would be around 72uW. Probably the current supply to the LED became too narrow in the Okutomi-kun's circuit (which was Akutsu's direction, sorry). The power supply of the driver is +/-14V, so even if the LED become brighter, the PD can deal with it.

Anyway, firstly I just setup the FFT analyzer (Agilent 35670A) so that it has DC coupling (the "range" was around 1.5V), but the dynamic range of the measurement system gets ruined due to the large DC voltage described above (see the graph attached).

Then today I simply set AC coupling and then the "range" can be reduced to around 12mV, and the measurements appear to be meaningful so far. As I'm not sure well what will happen in the lower frequency range when you select the AC coupling mode, so it would be better to use SR560 to cut DC signal because we can measure the transfer function and can be calbirated afterwards.

Images attached to this report
114_20150624173035_dcim1287.jpg 114_20150624173043_dcim1288.jpg 114_20150624175258_plot201506242.png 114_20150624175304_plot20150624ac.png
Comments related to this report
TomotadaAkutsu - 23:13, Friday 26 June 2015 (118)

[Abstract]

For measuring the LED intensity noise, the large DC voltage spoils the dynamic range of the FFT analyzer. The AC coupling could be usable to overcome the situation, so the difference between the AC and DC coupling inputs of 35670A's are measured. We should include the fit function for the calibration of measured data with the AC couping input.

[Measurements]

Measuring an input signal (or noise) with Ch1 and 2 of 35670A simultaneously and make it calculate the ratio Ch2/Ch1. A signal input is divided by a T-shape stuff and input into each channel.

  • (a) Freq. response of Ch2/CH1 are measured with a random noise input created by the 35670A itself.
  • (b) Swept sine mode to measure the same thing.

For the both cases, Ch2 is set to AC float and Ch1 to DC float. (I've tried other combination for just a confirmation.)

[Results]

The results of (a) and (b) are consistent, and  the fit transfer function is approximated by

f(x) = (i*x/x0)/(1+i*x/x0)

and

 x0 -> 0.485(5),

where I just assume f(x) -> 1 as x -> infinity.

You can use this function for the calbration when you use this 35670A with the AC coupling input (CH2).

Note: I'd like to avoid SR560s for the OSEM screening setup, as it appears to have large noise comparing to what I want to measure in the low frequency region, and each SR560 looks to have different noise level one by one... maybe our SR560s are getting old since they had been bought in 1998; one of them even sometimes has a jump in the DC offset during the measurement.

Images attached to this comment
118_20150626164136_bode20150626.png
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 19:03, Wednesday 24 June 2015 (113)Get code to link to this report
AC noise with Function Generator

I measured again the noise of the photodetector. The chopper is off and the lockin trigger comes from a function generator set at 430Hz. In order to check if the noise could come from the function generator, I measured the noise with the probe laser ON and with the probe laser OFF.

The AC noise std with the Probe ON is 18microV as usual.

The AC noise std with the Probe OFF is 0.7microV. Very similar to the case Chopper ON - Probe OFF measured before (0.8microV).

This means that the noise doesn't come from the function generator.

Other possible noise sources could be the HeNe power fluctuation or jitter position fluctuation due to some vibrations.

----------------------------------

During many noise acquisitions I noticed that sometimes (one time in 3 or 4 minutes) there is a strong peak of noise, about 2  order of magnitude higher than the usual noise. The system manual speaks about them and give this explanation: "There are various sources of spiky noise including dust particles crossing the probe beam path." 

General (Test)
Print this report.
TomotadaAkutsu - 10:35, Wednesday 24 June 2015 (112)Get code to link to this report
AEL: Satellite box preliminary test

A satellite box prototype was delivered to NAOJ last week, and was insitu revised/modified, having discussions with AEL group.

So the prototype circuit in NAOJ is not the final version; some capacitaces are removed due to their wrong assembly directions, D101 diodes are removed due to similar reason, and a Q101 transistor of Ch1 is removed during the investigation of the circuits.

Anyway, a basic measurement, the noise level of the output of the circuit is preliminarily measured. A PD (S1223-01) in a old design holder is connected to the circuit with a barrack conversion board, and the output's differential outputs are directly connected to a FFT analyzer (Agilent 35670A) with a front end setup of DC float.

The measurement are done during the room lights are turned off as much as possible, and the PD is covered by an easy carton box. So far, for some reason (probably large 50Hz signal), the FFT's range should be set no less than 158 mV, and the most of the measured data are buried under the measurement system noise.

Anyway, the vertical axis is converted to input equivalent current noise density; divided by 2 due to its differential output, and divided by 38.3k Ohm of the transimpedance; see http://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=3499

Note that this is not a test for OSEM but just a basic measurement of the circuit itself. To say something about OSEM, the output should be calibrated to the unit of "m/rtHz".

I'm wondering why the 50Hz apperas such large in the circuit; comparing to the measurement result of an oplev driver so far, it appears a little bit too large...

Images attached to this report
112_20150624033500_figure1.png
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 22:50, Tuesday 23 June 2015 (111)Get code to link to this report
Calibration history

Members: Manuel, D.Tatsumi.

We made the calibration of  bulk and surface reference samples, in three different days.

For the bulk calibration, we found a difference of about 2% among the days.

For the surface calibration, we found a difference of about 10% among the days.

The plots show the AC/DC ratio during the scan along the z axis.

Images attached to this report
111_20150623154754_calibhistorybulk.jpg 111_20150623154802_calibhistorysurf.jpg
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 00:25, Tuesday 23 June 2015 (109)Get code to link to this report
AC noise
Members: Manuel, D.Tatsumi
 
We measured the reference signal from the chopper: it is a square waveform of 5V p-p at a frequency of 430Hz
We set a waveform generator at 430Hz but the knob was too sensitive to touch and the best we could do was 426Hz (the lockin time constant is 100ms, so we can be inside ±10Hz).
We aquired 3 minutes of AC signal. Noise standard deviation is 19microV, wich is comparable with the noise measured with the real chopper (18microV).
This means that the chopper vibration do not dominate the noise.
Images attached to this report
109_20150623042052_chopnochop.jpg